### v1

# Closed-Form Interpretation of Neural Network Classifiers with Symbolic Regression Gradients

Abstract
I introduce a unified framework for interpreting neural network classifiers tailored toward automated scientific discovery. In contrast to neural network-based regression, for classification, it is in general impossible to find a one-to-one mapping from the neural network to a symbolic equation even if the neural network itself bases its classification on a quantity that can be written as a closed-form equation. In this paper, I embed a trained neural network into an equivalence class of classifying functions that base their decisions on the same quantity. I interpret neural networks by finding an intersection between this equivalence class and human-readable equations defined by the search space of symbolic regression. The approach is not limited to classifiers or full neural networks and can be applied to arbitrary neurons in hidden layers or latent spaces or to simplify the process of interpreting neural network regressors.
Keywords: Artificial Neural Networks, Symbolic Regression, Interpretation of Neural Networks
1 TL;DR
What do you do in practice? To interpret an artificial neural network, I fit the gradients of a symbolic regression model to the gradients of the output neuron of the neural network.
2 Introduction
Artificial neural networks have become the most powerful tool in machine learning applications in a wide range of applications such as image classification, speech recognition, or natural language processing. In many fields, artificial neural networks have achieved better than human performance, and the number of these fields is consistently growing. The larger and more powerful the artificial neural networks become the more elusive is their decision making to us humans. The black-box nature of these models makes it very challenging to comprehend their decision-making processes.
Interpreting artificial neural networks is a central endeavor in many applications. As artificial neural networks are entering domains where a detailed understanding of the decision-making process is crucial, it becomes increasingly important to interpret their inner workings. These fields contain safety-critical, medical [1, 2] applications where human lives could be in danger. The legislative domains [3, 4] in which each decision needs to be supported by law. Self-driving cars are in danger of causing accidents because of a lack of understanding of the underlying machine learning algorithms [5]. Further, in many scientific applications [6] the primary utility of a model is to help understand an underlying phenomenon much more than making accurate predictions.
There has been a lot of progress in identifying which input features contribute the most to a certain decision [7, 8, 9]. Using these approaches it is possible to identify where danger lingers in an image of a traffic situation or find malignant cancer cells in MRI scans. However, from this understanding of low-level features, it is still a long road to develop a comprehensive understanding of the concepts an artificial neural network learns in order to make its predictions. The review [10] provides an overview of explainable artificial intelligence, covering methods for understanding, visualizing, and interpreting deep learning models.
There are many examples, especially in artificial scientific discovery, where a low-level interpretation is not enough. In many scientific fields, it is imperative to gain a deep understanding of the concept that is underlying neural network predictions. Some of the most central scientific tasks are a) understanding the dynamics of systems b) explaining new phases of matter and c) finding conserved quantities and symmetry invariants. In recent years, many scientists have developed methods to approach these problems. There are methods for the automated discovery of the dynamics of a system [11], symbolic regression for conserved quantities [12], and support vector machines that map out phase diagrams [13, 14]. These methods are inherently interpretable. Similarly, much more powerful neural network approaches have been developed to address these problems, which tend to be more successful in learning the underlying concepts but conceal these concepts from a human scientist. However, there have been publications that report successful strategies to discover the underlying concepts from neural networks. These include understanding the equations behind a dynamical system [15, 16, 17], extracting phase signatures from a neural network [18, 19] or revealing conserved quantities and symmetry invariants with neural networks [20, 21, 22]. It is important to note that many of the above examples are tackled by neural network classifiers while most symbolic interpretation literature focuses on regression problems.
To address the problem of interpreting neural network classifiers, it is important to acknowledge that concepts encoded in an artificial neural network are stored in a highly convoluted and elusive manner. A neural network does not encode them in a human-readable form, but through complex interactions of thousands if not millions of neurons/perceptrons. In other words, even if the neural network learns a concept that could be put into a human-readable form, a neural network hides this concept through complex and highly nonlinear transformations.
This observation leads to the initial idea behind the interpretation framework proposed in this manuscript. Is it possible to extract a human-readable concept $g$ that is learned by a neural network $F$ by getting rid of the elusive transformation $\phi$ that conceals it.
$F(x_{1},...,x_{n})=\underbrace{\text{sigmoid}}_{\text{activationfunction}}(\underbrace{\phi}_{\text{uninterpretable transformation}}(\underbrace{g(x_{1},...,x_{n})}_{\text{closed form decision function}}))$ (1)
The current paper provides an answer to this question by embedding the neural network $F$ into an equivalence class of functions that learn the same concept. A closed-form expression of this concept can be recovered by finding the intersection of this equivalence class with the space of human-readable functions defined through a symbolic regression search space.
2.1 Outline
This article is structured in the following manner: After exploring the most closely related prior research on neural network interpretability in Section 3:Prior Work, I describe the formal and mathematical framework that is necessary for the interpretation algorithm in Section 4: Theoretical Framework. Next, I explain the computational implementation of the interpretation algorithm in Section 5:Interpretation Algorithm. Finally, I perform experiments demonstrating interpretation results in Section 6:Experiments. The findings and potential follow-up research are summarized in Section 7: Conclusion.
3 Prior Work
To the best of my knowledge, there are no successful approaches that manage to find closed-form expressions of the high-level concepts learned by neural network classifiers beyond reducing the problem to perceptrons.
The most closely related works that also embed the neural network into another function space are about symbolic metamodels[23, 24, 25], where the authors reproduce the output of an artificial neural network by minimizing the difference between the neural network output and so-called Meijer G-functions. However, this approach has so far only been successfully employed for low-dimensional regression problems.
On the other hand, it is possible to directly employ symbolic classification approaches to obtain a symbolic solution to a classification problem. This can be done by performing symbolic regression with margin or hinge losses.
Common symbolic regression libraries that mostly employ genetic algorithms include Eureqa, which was employed in [12], Operon C++[26], PySINDy [27], Feyn[28], Gene-pool Optimal Mixing Evolutionary Algorithm [29], GPLearn [30] and PySR[31]. There are also symbolic regression algorithms that employ neural networks, for example, EQL with embedded symbolic layers[32, 33], deep symbolic regression uses recurrent neural networks [34], and symbolic regression with transformers[35]. Further, AI Feynman uses neural networks to simplify expressions [36]. An overview of interpretable scientific discovery with symbolic Regression can be found in[37]
4 Theoretical Framework
Figure 1: An artificial neural network is a connected graph consisting of nodes representing neurons and weighted connections between them. A neural network predicts an approximate target $F(\mathbf{x})=\hat{y}\approx y$ here in the context of binary classification. Removing the final sigmoid activation function allows the extraction of the latent model $f$ from the full neural network $F$ for easier interpretability.
4.1 Artificial Neural Network
An artificial neural network fig. 1 is a connected graph containing nodes representing neurons otherwise known as perceptrons and connections representing weighted inputs that are supplied to these nodes. A basic feedforward neural network can be represented mathematically as follows:
A neural network $F$ is applied to an input data point $\mathbf{a}^{(0)}=\mathbf{x}\in\mathbb{R}^{n}$ represented through an embedding in an $n$-dimensional space of real numbers. The input gets processed through a number $l$ of layers $L$ containing weights matrices $\mathbf{W}\in\mathbb{R}^{m_{l}\times m_{l-1}}$, biases vectors $\mathbf{b}\in\mathbb{R}^{m_{l}}$ and non-linear activation functions $\sigma$
$\mathbf{a}^{(l)}=L^{l}(\mathbf{a}^{(l-1)})=\sigma^{(l)}(\mathbf{W}^{(l)}\mathbf{a}^{(l-1)}+\mathbf{b}^{(l)})$ (2)
to produce some output
$\mathbf{\hat{y}}=F(\mathbf{x})=(L^{(l)}\circ...\circ L^{(0)})(\mathbf{x})$ (3)
The parameters of an artificial neural network are trained on a data set $(X,Y)$ representing the data manifold $X\subset D\subset\mathbb{R}^{n}$ through gradient descent/backpropagation to minimize an objective/loss/error function $L(F(X),Y)$ such that $F(X)\approx Y$ on all data points on the data manifold.
4.2 Interpreting an Artificial Neural Network
Interpreting an artificial neural network fig. 1 means formulating a mapping between an abstract mechanism or encoding into a domain that a human can understand. Human understanding is a vaguely defined concept since human capabilities range from simply perceiving common objects in the real world to expert knowledge in their respective field. I suggest categorizing interpretations along the following four different attributes, which is more fine-grained than the categories proposed in [10]:
A) Mechanistic vs. Functional: Mechanistic interpretation is concerned with explaining the mechanisms the neural network employs to solve a problem or to implement a function. Functional interpretation is about understanding the concepts by which a neural network relates the input to the output.
B) Local vs. Global: Local interpretation is about understanding which elements of a certain data point influence the prediction, while global interpretation explains what features are relevant for a learned concept in general.
C) Verifying vs. Discover: A machine learning practitioner might have a set of hypotheses of concepts in mind that a neural network might learn in order to solve a given task. It is often easy to verify or falsify if these concepts are included in the features learned by an artificial neural network. However, discovering an unknown concept without a list of potential hypotheses is in general very hard.
D) Low-Level vs. High-Level Features: The neurons of the network jointly implement a complex nonlinear mapping from the input to the output through potentially several hidden layers. The global concept that must be interpreted is usually represented by a neuron in the final layer. Low-level features that are close to the input are usually rather interpretable while concepts in the final layers or some latent space are in general very abstract.
Interpreting an artificial neural network through the framework presented in this paper falls into the category of functional global discovering high-level interpretation.
Figure 2: The interpretation framework is based on finding the intersection between human-readable functions and the equivalence class (6) of functions that contain the same information as an output neuron of the neural network. The space of human-readable functions can be defined through a symbolic regression search space with elementary functions and complexity that matches the user’s knowledge.
4.3 Equivalence Class of Functions Containing the Same Information
4.3.1 Definition
The goal of the interpretation framework outlined in this manuscript is to find the intersection of the equivalence class of functions that contain the same information as the output neuron (or any other neuron) and the set of human-readable equations, see fig. 2.
For this purpose, it is important to define what it means that a certain neuron in an artificial neural network learns to encode a certain quantity. This neuron can be seen as a function $f(\mathbf{x})$ that depends on the input $\mathbf{x}$. $f(\mathbf{x})$ contains the full information about a certain quantity $g(\mathbf{x})$ if $g(\mathbf{x})$ can be faithfully reconstructed from $f(\mathbf{x})$. Conversely, if $f(\mathbf{x})$ only contains information from $g(\mathbf{x})$ it is possible to reconstruct $f(\mathbf{x})$ from the knowledge of $g(\mathbf{x})$. In mathematical terms that means that there exists an invertible function $\phi$ such that $f(\mathbf{x})=\phi(g(\mathbf{x}))$.
For the purpose of this work, $g(\mathbf{x}),f(\mathbf{x})\in C^{1}(D\subset\mathbb{R}^{n},\mathbb{R})$, are assumed to be continuously differentiable functions, where $C^{1}(D\subset\mathbb{R}^{n},\mathbb{R})$ denotes the vector space of differentiable functions from $D$ to $\mathbb{R}$. Further, $D\subset\mathbb{R}^{n}$ is the data manifold which is required to be compact and simply connected. By employing these definitions it is possible to define an equivalence set of all functions that extract the same scalar-valued information from the data:
$\tilde{H}_{g}=\left\{f(\mathbf{x})\in C^{1}(D\subset\mathbb{R}^{n},\mathbb{R})\ |\exists\text{ invertible }\phi\in C^{1}(\mathbb{R},\mathbb{R})\ :\ f(\mathbf{x})=\phi(g(\mathbf{x}))\right\}$ (4)
If $f(\mathbf{x})$ is the output of an artificial neural network, this defines an equivalence relation between all functions that base their decision on the same quantity $g(\mathbf{x})$. Several neural networks might encode the same function $f$ because it does not matter how the output is calculated mechanistically. All realizations of a neural network that is symmetric under permutations or other changes in its weights would still be functionally equivalent and thus the same function within the equivalence class $\tilde{H}_{g}$.
Since it is computationally very inefficient to verify whether two functions or neural networks belong to the equivalence class $\tilde{H}_{g}$, according to the above definition, it is convenient to reformulate it. Taking the gradient of a function $f\in\tilde{H}_{g}$ with respect to the input yields
$\nabla f(\mathbf{x})=\phi^{\prime}(g(\mathbf{x}))\nabla g(\mathbf{x})$ (5)
Since $\phi$ is invertible, $\phi^{\prime}$ is strictly monotonic, thus $|\phi^{\prime}(g(\mathbf{x}))|>0$, hence the gradients of $f(\mathbf{x})$ and $g(\mathbf{x})$ are always parallel $\nabla f(\mathbf{x})||\nabla g(\mathbf{x})$ at every $\mathbf{x}\in D$.
Using this property, I define the equivalence set $H_{g}\subseteq\tilde{H}_{g}$, which will later form the basis for the neural network interpretation algorithm:
$H_{g}=\left\{f(\mathbf{x})\in C^{1}(D\subset\mathbb{R}^{n},\mathbb{R})\ |\ \frac{\nabla f(\mathbf{x})}{\lVert\nabla f(\mathbf{x})\rVert}=\frac{\nabla g(\mathbf{x})}{\lVert\nabla g(\mathbf{x})\rVert}\lor\nabla f(\mathbf{x})=\nabla g(\mathbf{x})=0,\ \forall\mathbf{x}\in D\right\}$ (6)
[…]
It remains to show that by using the above definitions: $H_{g}=\tilde{H}_{g}$. Hence, one needs to show that for each function $f\in H_{g}\ \exists\phi:f(\mathbf{x})=\phi(g(\mathbf{x}))$. It is possible to explicitly construct the function $\phi$ that maps between $f$ and $g$. Defining $\phi^{\prime}$ through (5) leads to an integrable $\phi^{\prime}(g(\mathbf{x}))\nabla g(\mathbf{x})=\nabla f(\mathbf{x})$ because a) the images of $f(D)$ and $g(D)$ are compact, thus $\phi^{\prime}$ maps between compact subsets of $\mathbb{R}$ and b) $\phi^{\prime}$ is continuous. For any simply connected $D\subset\mathbb{R}^{n}$ I can define the $C^{1}$-curve $\mathbf{x}:[t_{0},t_{1}]\rightarrow D$.
[…]
$=f(\mathbf{x}(t_{1}))-f(\mathbf{x}(t_{0}))$ (11)
Note, since I base the integral on the product $\phi^{\prime}(g(\mathbf{x}))\nabla g(\mathbf{x})$ instead of $\phi^{\prime}(g(\mathbf{x}))=\lVert\nabla f(\mathbf{x})\rVert/\lVert\nabla g(\mathbf{x})\rVert$, I avoid problems arising from $\nabla f(\mathbf{x})=\nabla g(\mathbf{x})=0$. Similarly, one can proof the existence of $\tilde{\phi}:\tilde{\phi}(f(\mathbf{x}))=g(\mathbf{x})$ such that $f(\mathbf{x})=\phi(\tilde{\phi}(f(\mathbf{x})))$ and thus $\phi$ is invertible. Having explicitly constructed $\phi$ proofs $H_{g}=\tilde{H}_{g}$.
4.3.2 Assumptions
In a practical machine learning application not all assumptions from the prior section that ensure $H_{g}=\tilde{H}_{g}$ hold true. However, even then $H_{g}\approx\tilde{H}_{g}$ can provide a good approximation that allows for a retrieval of the function that a neuron encodes.
A machine learning data set $D\subset\mathbb{R}^{n}$ is bounded since it is finite. However, if there is a divergence in the function that the machine learning model is supposed to approximate, the data set might not be closed and thus not compact. A data set $D$ might not be simply connected, especially if it is in the form of categorical data or images.
A neural network classifier, if successfully trained, tends to approximate a categorical output, which is neither continuous nor differentiable. However, this binary output is typically an approximation mediated by sigmoid or softmax activation functions, which indeed are continuously differentiable. Still, interpreting artificial neural networks with the framework introduced in this paper experiences numerical artifacts if a gradient is taken from a network that contains sigmoid or softmax activation functions. For this reason, I suggest avoiding these activation functions in the design of hidden layers and removing them from the output neuron during the interpretation process (the same argument holds true for tanh or related activation functions).
The above definitions of equivalence classes could be extended to piecewise $C^{1}(\mathbb{R}^{n},\mathbb{R})$ functions. This function set contains many artificial neural networks that include piecewise differentiable activation functions like $\text{ReLU}(x)=\max(0,x)$. However, this causes problems when evaluating derivatives close to ReLU ($\text{ReLU }(x)=0$. In practice, one can observe that piecewise $C^{1}$ activation functions lead to computational artifacts when calculating gradients. Hence, I suggest using $\text{ELU}=\exp(x)-1|x\leq 0,x|x>0$ as the preferred activation function in hidden layers.
4.4 Symbolic Regression Search Space
4.4.1 Human-Readable Functions
An interpretable representation of a neural network is a description that humans can make sense of. Each single neuron can be described as a simple linear mapping together with an activation function that can be expressed in terms of a short equation. However, the complexity that arises from combining several thousand or even millions of these neurons conceals any meaning of the function that is approximated by a neural network. In order to discuss the requirements for interpretable representations in the language of mathematics, let me discuss three criteria that are important for comparing representations of equations.
1)Building blocks: In the realm of mathematical equations human readable equations are typically written in terms of numbers (integers, real numbers, complex numbers), binary ($+$,--,$\times$) operations, or unary operations that summarize complex operations or arise from the solutions of ubiquitous equations ($d/dx$,$\int\cdot\ dx$,$\exp(x)$,$\sin(x)$,$x!$).
2) Complexity: Humans can understand short combinations of the above elementary constituents of equations, however, combining a large number of them prevents interpretability. Hence, to formulate complex objects, it is helpful to summarize and simplify repeating processes in equations to a shorter form. For example, the Lagrangian of the standard model of particle physics [38] fits on a coffee mug in its most simple formulation. However, each element in the equation represents a collection of more complex objects that are conveniently grouped together.
3) Context: Pricing financial derivatives with the Black-Scholes equation [39] requires a firm understanding of the assumptions made to derive it. Identifying something as a Lagrangian or a conserved quantity means associating a certain functional purpose or properties to an equation. For example, reading the standard model Lagrangian in its full glory is a futile exercise, unless the reader has a deep understanding of what role Lagrangians play in the larger framework of quantum field theory. The same can be said about much simpler equations like the de-Broglie wavelength $\lambda=h/p$ ($h$-Panck constant, $p$-momentum) [40], which is arguably one of the simplest physics equations. This equation is meaningful because of the context of its formulation, it is instrumental in associating a quantum mechanical wavelength to massive objects.
Throughout this paper, I base the set of human-readable functions on symbolic regression trees, see fig. 3. These trees can be constructed to contain a user-defined set of elementary operations. Symbolic trees can also be associated with a complexity measure that increases the larger the tree grows leading to several solutions along the Pareto front. Hence, it is possible to address human readability according to aspects 1) and 2). Within the scope of this paper, it is impossible to address the question of the context that gives the equation meaning, as discussed in 3). In recent years, progress in artificial scientific discovery has made it possible to train neural networks to optimize loss functions based on concepts instead of regression targets [18, 20, 22, 21]. The interpretation framework developed in the current manuscript is uniquely suitable to address the interpretation in these cases.
Figure 3: Symbolic regression tree that represents a mathematical function. A tree consists of connected nodes containing variables, numeric parameters, unary and binary operators. Symbolic regression is performed by a genetic algorithm that modifies, evolves and adds nodes to optimize an objective function on some underlying training data.
4.4.2 Symbolic Regression Algorithm
In order to computationally search through the space of human-readable functions and determine the intersection with the equivalence set described in the previous section, it is important to build upon a suitable symbolic regression algorithm.
In principle, a simple way of searching through the space of possible functions could be optimizing a linear combination of user-defined elementary functions. A much better approach can be taken by utilizing evolutionary or genetic algorithms. These algorithms build trees of connected nodes that represent a mathematical function, see fig. 3. Each node represents input variables, numeric parameters, as well as unary and binary operators. A genetic algorithm modifies, evolves, and adds nodes to optimize an objective function on some underlying training data.
Many such algorithms have been developed in recent years, these include: Eureqa [12], Operon C++ [26], PySINDy [27], Feyn [28], Gene-pool Optimal Mixing Evolutionary Algorithm [29] and GPLearn [30]. Other symbolic regression techniques based on neural networks are EQL [32, 33], AI Feynman [36], Deep symbolic regression [34] and Symbolic regression with transformers [35].
In this paper, I build upon PySR [31], a high-performance symbolic regression algorithm for Python. PySR’s internal search algorithm is a multi-population genetic algorithm, which consists of an evolve-simplify-optimize loop, designed for the optimization of unknown scalar constants in newly-discovered empirical expressions. In order to optimize the computation speed, PySR’s backend is coded in Julia as the library SymbolicRegression.jl.
5 Interpretation Algorithm
In this section, I discuss how to interpret a neural network classifier $F$ in practice by finding a representative symbolic function that lives on the intersection between human-readable functions and all functions that contain the same information as a trained neural network, see fig. 2. While the description here focuses on binary classifiers, it is possible to apply the framework to interpret multi-class classifiers or even neurons contained in the hidden layers of neural networks. The interpretation framework contains three algorithmic steps: 1) training an artificial neural network 1, 2) extracting gradients of the output neuron with respect to the input 2 and 3) performing symbolic regression to find a symbolic function whose gradients approximate the gradients from the desired neuron of an artificial neural network 3.
Figure 4: (a) Two class data of Experiment 1 separated by decision boundary $g(\mathbf{x})=x_{1}^{2}+2x_{2}^{2}=1$. A neural network $F$ is trained to classify the data. Afterward, a symbolic model $T$ is trained to reproduce the normalized gradients of $F$ which coincide with the normalized gradients of function that defines the decision boundary $g$. (b) Empirical correlation between true function $g$ and the neural network $F$. Removing the sigmoid activation function from $F$ defines the latent model $f$ which has an almost linear correlation with $g$. However, this correlation is not linear and defines the function $f=\phi(g)$ with which I ascertain the equivalence relation $f\sim g$ assuring that $F$ and $f$ contain the same information as $g$ and thus $F,f\in\tilde{H}_{g}$ according to eq.(4).
5.1 Training a Neural Network For Binary Classification
The first part of the interpretation procedure involves training an artificial neural network for binary classification on a two-class data set, see algorithm 1, to approximate the decision boundary to separate the two classes, see fig. 4(a). After successful training I obtain the full model $F$ and the latent model $f$ by removing the final sigmoid activation function, which has a cleaner correlation to a closed-form expression of the decision boundary $g$, see fig. 4(b):
$F(\mathbf{x})=\text{sigmoid}(f(\mathbf{x}))$ (12)
I implemented one neural network architecture for all experiments using the Python library tensorflow [41]. This neural network consists of two hidden layers with 1000 neurons each and ELU activation functions. I compared different activation functions and ELU activations yielded gradients least affected by computational artifacts. The final layer contains one single neuron without a sigmoid activation function, commonly used for binary classification. The weights and biases in the hidden layers are regularized with an $L^{2}$ penalty of $10^{-4}$. Further, I employ dropout regularization after each hidden layer with a chance of $20\%$. I train the neural networks using the Adam optimizer and use learning rate and early stop callbacks that reduce the learning rate by 50% or stop training if the loss stops decreasing. For this reason, it is enough to set the number of epochs large enough such that the early stopping is always triggered. The batch size in all cases is 100. I train on $80\%$ of the data set and use another 20% for validation. Since I am not interested in calculating any test accuracy here, the validation and test set are the same.
Data: Labelled data set $D=(X_{train},Y_{train})$
Input: Neural Network Hyperparameters
1 Initialize neural network classifier with sigmoid activation at output $F$
2 Train $F$ on $D$
$f$leftarrow remove sigmoid activation of output neuron of $F$ Output: Trained model $F$, latent model $f$
Algorithm 1 Train Neural Network for Binary Classification
Data: Labelled data set $D=(X_{train},Y_{train})$
Data: Unlabelled data set $D_{u}=(X_{u})$triangleright Artificial data not used for training
Input: Trained model $F$ and latent model $f$
Data: Selection threshold $\delta$
1 $\tilde{X}\leftarrow(X_{train},X_{u})$triangleright Add artificial unlabelled data
2 $\tilde{X}\leftarrow\tilde{X}\textbf{ where }F(\tilde{X})\in[\delta,1-\delta]$triangleright Select data close to decision boundary
3 $G_{f}\leftarrow[\nabla f(\mathbf{x})\text{ for }\mathbf{x}\text{ in }\tilde{X}]$triangleright Gradients of latent model wrt. input
4 $G_{f}\leftarrow[\textbf{if }\nabla f(\mathbf{x})\neq 0:\nabla f(\mathbf{x})/\lVert\nabla f(\mathbf{x})\rVert$
[…]
Output: $(\tilde{X},G_{f})$
Algorithm 2 Obtain Gradients of Latent Model
Data: Gradient data set $(\tilde{X},G_{f})$
Input: Symbolic Regression Hyperparameters
Data: Set of unary and binary operations.
1 initialize symbolic regression model $T$
2 evolve $T$ with (
3 aaa $G_{T}\leftarrow[\nabla T(\mathbf{x})\text{ for }\mathbf{x}\text{ in }\tilde{X}]$triangleright Gradients of symbolic model
4 aaa $G_{T}\leftarrow[\textbf{if }\nabla T(\mathbf{x})\neq 0:\nabla T(\mathbf{x})/\lVert\nabla T(\mathbf{x})\rVert$
[…]
6 aaa to minimize MSE$(G_{f},G_{T})$
Output: Symbolic regression model $T$
Algorithm 3 Symbolic Search
5.2 Obtaining Gradients of Latent Model
The second step is described in algorithm 2, which is used to collect the gradient information, see fig. 4(a), for the symbolic regression step. It involves potentially adding additional unlabelled data points to the training data set $X$, either from available unlabelled data, sampling from the data manifold, or by perturbing training data. This data set is denoted $\tilde{X}$. I omitted increasing the training data set in my experiments, but in data-scarce domains, this could lead to a significant improvement in the symbolic regression results.
[…]
Afterward, I calculate the normalized gradients $G_{f}=\nabla f/\lVert\nabla f\rVert$ of the latent model $f$, with respect to the input variables, on the remaining data points. Finally, together with the inputs they get stored in the form of a labeled data set $(\tilde{X},G_{f})$ that is used as training data of the symbolic regression algorithm.
5.3 Symbolic Search
The third step of the interpretation algorithm uses symbolic regression, see fig. 3, to find human-readable equations describing the decision boundary $g$. The algorithm is summarized in algorithm 3. This step employs the backend of the Python symbolic regression module PySR [31] coded in Julia called SymbolicRegression.jl.
The interpretation framework proposed in this paper has requirements beyond what symbolic regression algorithms typically do. Symbolic regression models are represented by trees $T$ that are trained on a labeled data set $(X,Y)$ to reproduce an output $Y$. However, the task in this manuscript is to perform symbolic regression on the normalized gradient $G_{f}=\nabla f/\lVert\nabla f\rVert$ of a specific neuron within the neural network. Thus, the gradient set $G_{f}$ takes on the role of the data set label $Y$. Further, the symbolic regression tree must be differentiable in order to calculate the normalized gradients of the tree $G_{T}=\nabla T/\lVert\nabla T\rVert$ itself. Finally, the desired solution of the symbolic regression procedure is not finding a closed form expression of the normalized gradient $\nabla f/\lVert\nabla f\rVert$, but the function $f$ itself. An additional limitation of SymbolicRegression.jl is its limitation to one-dimensional regression targets in the standard form.
The extremely customizable library SymbolicRegression.jl can be programmed to 1) accept gradient information as data set, 2) calculate gradients of a symbolic regression tree 3) normalize these gradients and 4) optimize a custom loss function that compares normalized gradients.
I choose the mean square error (MSE) loss function between latent model normalized gradient $G_{f}$ and normalized gradient of the symbolic tree $G_{t}$ as the most straight-forward choice of the objective function to ascertain that both $f$ and $T$ fall into the same equivalence class defined by eq.(6). This loss function is equivalent to the cosine loss (CSL) based on the scalar product of two vectors.
$\text{MSE}(G_{f},G_{T})$ $=\frac{1}{n}\sum_{i=1}^{n}\left(\lVert G_{f,i}-G_{T,i}\rVert^{2}\right)$ (13)
$=\frac{1}{n}\sum_{i=1}^{n}\left(\underbrace{\lVert G_{f,i}\rVert^{2}}_{1}+\underbrace{\lVert G_{T,i}\rVert^{2}}_{1}-2G_{f,i}\cdot G_{T,i}\right)$ (14)
$=2-2\text{CSL}(G_{f},G_{T})$ (15)
The symbolic regression model is trained to find a closed-form function corresponding to the decision boundary using the following building blocks:
  1. 1.
    input variables $x_{1},...,x_{n}$
  2. 2.
    floating-point constants
  3. 3.
    binary operators: +, -, *, /
  4. 4.
    unary operators: sin, exp
The hyperparameters are a batch size of 25, a maximum number of epochs of 200, an early stop callback, and a maximum tree size of 30 nodes.
The output of the symbolic regression algorithm is not just a single closed-form function, but a set of symbolic functions that balance complexity and accuracy along the Pareto front. More precisely, the result contains functions with low complexity and low accuracy, and more complex functions are added if the complexity can be justified by an increase in accuracy.
Table 1: Data sets
Variables Decision Formula
Experiment 1 2 $x_{1}^{2}+2x_{2}^{2}=1$
Experiment 2 2 $x_{1}^{2}+3x_{1}x_{2}+2x_{2}^{2}=5$
Experiment 3 3 $x_{1}^{2}+\sin(x_{2}+x_{3})=1$
Experiment 4 3 $\exp(x_{1}-x_{2})-\pi x_{3}=1$
Experiment 5 3 $x_{1}\exp(-\frac{1}{2}(x_{2}^{2}+x_{3}^{2}))=\frac{1}{5}$
Experiment 6 6 $\frac{x_{1}x_{2}}{(x_{3}-x_{4})^{2}+(x_{5}-x_{6})^{2}}=1$
6 Experiments
6.1 Data
I apply the outlined interpretation procedure to several neural networks that are tasked with classifying data sets that are each grouped into two classes. The classes are separated through a domain boundary defined by a decision formula $g$. The interpretation framework attempts to recover these functions from the neural networks. I perform experiments on six different data sets outlined in Table 1. For each data set, I sample 10000 data points and categorize them into classes according to the decision formulas. In all cases, the data sets are created with multiplicative Gaussian noise of $1\%$.
Figure 5: The results of fitting a symbolic regression model $T$ to the normalized gradients of the neural network are presented along the Pareto front. The Pareto front collects several possible results with decreasing Mean Square Error (MSE) and increasing complexity. The closest match to the true underlying function is often found at the point of steepest change of the Pareto front.
6.2 Results
Applying the proposed interpretation framework to the six data sets yields sets of formulas along the Pareto front. All formulas are equally valid and trade complexity for accuracy. The Pareto fronts for the six experiments are collected in fig. 5.
The interpretation in experiment 1 is able to correctly recover the decision function $x_{1}^{2}+2x_{2}^{2}$. In experiment 2, the closest match is $-0.712x_{1}+x_{2}$ which seems incorrect at first glance. However, if we promote the transformation $\phi\rightarrow(\phi/0.712)^{2}$, we can map the current solution to a very good approximation of the true function $x_{1}^{2}+2.81x_{1}x_{2}+1.97x_{2}^{2}$. In experiment 3 the algorithm correctly recovers $x_{1}^{2}+\sin(x_{2}+x_{3})$. Similarly, promoting the transformation $\phi\rightarrow(\phi/0.32)^{2}$ in experiment 4 yields a very good approximation to the true function $\exp(x_{1}-x_{2})-3.125x_{3}$. In experiment 5, I recover the equation $x_{1}\exp(-0.479(x_{2}^{2}+x_{3}^{2}))$. However, in this experiment, the correct function is in a very flat region of the Pareto front. Lastly, experiment 6 fails to learn the division and instead replaces it with a subtraction $x_{1}x_{2}-((x_{3}-x_{4})^{2}+(x_{5}-x_{6})^{2})$. Most likely the algorithm learns a Taylor approximation $1/(1+x)\approx 1-x$ of the true function and removes any remaining constants since they do not contribute when calculating gradients. It is of note that in most experiments, the correct function can be found when the Pareto front exhibits the steepest change.
7 Conclusion
In this paper, I introduce a comprehensive framework to interpret neural network classifiers that is equally capable of interpreting single neurons in hidden layers or latent spaces. The solution presents the learned concept as a closed-form function. The interpretation method is based on the idea of finding an intersection between the space of human-readable functions and all functions that contain the same information as a specific neuron in an artificial neural network. Through the interpretation procedure, it is possible to get rid of uninterpretable transformations that conceal the true function and focus on the simplest form of the underlying concept.
I demonstrate the power of the interpretation method in section 6 at the example of six different data sets with two to six input variables each. In 4 cases the interpretation procedure recovers the exact functional form of the true decision function, while in 2 experiments the procedure finds a very good symbolic approximation. Comparing these results with other works is impossible since most related works only interpret neural network regressors. However, the most similar research articles manage to only solve problems based on one variable [23] or two variables variables [24].
The interpretation method can also be used to simplify the interpretation of neural network regressors. Consider for example
$G(x_{1},x_{2})=$ $\frac{a}{\exp(b\ x_{1}+c\sin(dx_{2}))+1}=\frac{a}{\exp(bz)+1}=g(z)\ ,$ (16)
with $\text{with }z=x_{1}+c/b\sin(dx_{2})$ (17)
To find a symbolic expression of a neural network learning the function $G(x_{1},x_{2})$ in the context of regression, use the framework in this paper to find $z$, then prepare the symbolic regression problem for $g(z)$ and use your favorite symbolic regression algorithm to solve a simple one-dimensional symbolic regression problem.
In the future, it might be worth exploring how well a neural network classifier + interpretation approach compares to directly applying symbolic classification methods. I suspect that a neural network is a superior learner and the artificial data generation in algorithm 2 enables a better symbolic performance. I also believe that the neural network can deal with possible data noise better such that the artificially created data for the symbolic regression steps could be essentially noise-free.
There is of course the limitation of not being able to learn functions of a single variable since the interpretation method’s goal is to reduce the concept learned by a neural network to a one-dimensional function. This also means one loses the numerical value of the threshold/bias of the output neuron. This is however not a problem, since solving a one-dimensional symbolic regression problem or fitting a threshold/bias is minor compared to the complexity of problems solved by the presented interpretation framework.
It remains a central open question: why does a neuron contain the concept when it evaluates the function far away from the decision boundary? In my intuition, there is only a very weak training signal induced by inputs that are already classified correctly.
When applying the interpretation algorithm to practical problems it might be useful to sort out equations from the Pareto Front that violate dimensional constraints. Similarly, dimensional analysis could also be used inside the symbolic regression algorithm itself.
Further, one might have some inductive bias in mind that is better represented through other function sets. There is no need to use symbolic regression trees to represent the solution, any suitable differentiable function space can do.
The code used for this project can be found at https://github.com/sjwetzel/PublicSymbolicNNInterpretation
8 Acknowledgements
I thank the National Research Council of Canada for their partnership with Perimeter on the PIQuIL. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development Canada and by the Province of Ontario through the Ministry of Economic Development, Job Creation and Trade. This work was supported by Mitacs through the Mitacs Accelerate program. I also thank the PySR team, especially Miles Cranmer, for their helpful comments on github.